# Wiedemann-Franz law violation domain for graphene and nonrelativistic systems

### v2

@@ IV Summary and Interpretation
In summary, we have gone through comparative research on Lorenz ratios for 3D, 2D NR and G systems for a systematic search of the WF law violation domain. During our analysis, we found the WF law violation domain of Fermi energy (normalized by a fixed temperature), where the Lorenz ratio deviates from one. First, we have used standard solid state book-based non-fluid expressions of Lorenz ratio for the 3D-NR case, which are also applied for 2D-NR, 3D-G, and 2D-G cases. For all cases, the Lorenz ratio remains constant in the higher values of normalized Fermi energy. This fact is well in agreement with the universal validity of the WF law in metal systems, which can be described by the 3D-NR case of the electron gas having larger Fermi energy ($2-10$ eV) or normalized Fermi energy 400-2000 at $60^{\circ}$ K. Depending on the definition of specific heat at constant volume and Fermi energy, we will get two possible saturated values of Lorenz ratios - 1 and 2, where former is experimentally observed (average) value but one may mark the saturating domain as the WF law obeying domain. If one identifies two domains - less and greater than $10$ value of normalized Fermi energy (i.e. $\frac{\epsilon_{F}}{T}<10$ and $\frac{\epsilon_{F}}{T}>10$), then the former domain of all cases show the deviation of WF law. This tuning of normalized Fermi energy is conveniently possible for 2D-G system only via tuning the electron doping. Recently measured Lorenz ratio by Crossno et al. [15] indicates the violation of the WF law in the lower normalized Fermi energy domain, which is popularly called as Dirac fluid (DF) domain. Using standard solid state or non-fluid theory for 2D-G system, we also find WF law violation in the lower normalized Fermi energy domain. However, non-fluid theory gets less than one values of Lorenz ratio, while experiment gets its greater than one values. It demands that we may have to transit towards a fluid-type description from the standard non-fluid expressions of the Lorenz ratio. So far, to the best of our knowledge, this straightforward thermodynamical phase space analysis of Lorenz ratio, based on non-fluid theory is missing in the literature. Present work is fulfilling this gap, which may be considered one of the important contents of present article. A next level content of present article is its fluid based calculations, whose outcomes are summarized in the next paragraph.
@@
Towards the fluid description, we have approached in two steps. In the first step, we have used Boltzmann transport equation for 2D-G system, where local thermalization concept is mainly implimented for fluid description. If we compare the NF and F expressions of Lorenz ratios, then we can find that only via specific heat, Fermi integral functions are entered into the NF expression of Lorenz ratio. On the other hand, F expressions of electric and thermal conductivity and hence Lorenz ratio carry additional Fermi integral functions, which are completely appeared because of local thermalization aspect of fluid description. During this first step attempt of fluid description, we get a little deviated outcomes of Lorenz ratio with respect to NF results. Sommerfeld limit of F-based Lorenz ratios give a divergent outcomes instead of constant values (it was found in its NF-based expression). This divergent trend of Lorenz ratio in lower normalized Fermi energy is experimentally observed, so this first step fluid description results may be considered as a preliminary hints of NF to F transition in the Dirac fluid domain or lower Fermi energy domain. Observing the quantitative mismatch between experimental data and first step fluid theory, we have proceeded to second step extension of fluid theory. Here, we have gone through the replacement of specific heat by enthalpy per net electron (divided by temperature). This replacement is based on the intuition for transition from quasi-particle concept to fluid concept. In NF picture, electron average energy is the quasi-particle concept and its temperature derivative gives specific heat concept. Whereas, enthalpy per net electron is the fluid concept for electron-hole plasma in Dirac fluid domain. Interestingly, we notice that this conceptual transformation and replacement provide an enhancement of Lorenz ratio in Dirac fluid domain and their theoretical values are quite close to experimental data points.
@@
After developing this step by step NF to F frameworks, we have identified four regions along the axis of normalized Fermi energy -
@@
  • Above $10$: Lorenz ratio is one and it can be called hardcore NF Ohmic domain. Sommerfeld limit of our NF expression can explain it easily.
@@
  • Within $2.5$ to $10$: Lorenz raio become less than one and it can be called NF but non-ohmic type domain. Our NF expression of Lorenz ratio and experimental data both hint this domain.
@@
  • Within $0.7$ to $2.5$: It is the transition domain from NF to F by showing less than one to greater than one values of Lorenz ratio.
@@
  • Below $0.7$: Here, Lorenz ratio becomes greater than one and it can be called Fluid domain as enhancing trend in fluid theory and experimental data towards lower doping range is noticed.
@@
At the end, for final quantitative data matching, we have demonstrated a sandwich estimation of F and NF results. However, present article intend to draw the attention of reader on the qualitative message of NF to F transition. This association of enhancement Lorenz ratio observation in the Dirac fluid domain with the NF to F transition is agreement with earlier theoretical works. Although, some alternative explanations are also addressed by some earlier works, more theoretical studies are probably required for better understanding about this WF law violation in graphene system.
@@ Appendix A FERMI-DIRAC DISTRIBUTION FUNCTION
$-\frac{\partial f_{0}}{\partial\epsilon}=\frac{\beta e^{\beta\left(\epsilon-\epsilon_{F}\right)}}{\left(e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1\right)^{2}}=\frac{\partial}{\partial\epsilon_{F}}\left(\frac{1}{e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1}\right).$ (80)
@@ Appendix B FERMI-DIRAC FUNCTION
    Case.2. When $A$ is too much small $\left(\epsilon_{F}<<k_{B}T\right)$, then the function becomes simplified as $A$
@@ Appendix C ELECTRONIC SPECIFIC HEAT
[…]
@@
[…]
@@
[…]
@@
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=\Bigg{[}\frac{\partial}{\partial T}\Big{(}\frac{U}{N}\Big{)}\Bigg{]}_{S,\ \epsilon_{F}}.$ (98)
@@
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg{[}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}+T\frac{\partial}{\partial T}\left(\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\right)\Bigg{]}.$ (99)
@@
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}~{}.$ (101)
@@
[$\prescript{2D}{G2}{[C_{V}]_{e2}}=\frac{1}{N}\frac{\partial U}{\partial T}\Bigg{|}_{S,\ \epsilon_{F}}.$ (102)
@@
[$\prescript{2D}{G2}{[C_{V}]_{e2}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (105)
@@
[$\prescript{2D}{NR1}{[C_{V}]_{e1}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]},$ (106)
@@
[$\prescript{2D}{NR2}{[C_{V}]_{e2}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (107)
@@
[$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]},$ (108)
[$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (109)
[$\prescript{3D}{G1}{[C_{V}]_{e1}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]},$ (110)
[$\prescript{3D}{G2}{[C_{V}]_{e2}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (111)

### v3

@@ IV Brief Interpretation of Results
In this section, we will quickly revisit our results and try to interpret them briefly in terms of the transformation from non-fluid to fluid description. Here, we have gone through comparative research on Lorenz ratios for 3D, 2D NR and G systems for a systematic search of the WF law violation domain. During our analysis, we found the WF law violation domain of Fermi energy (normalized by a fixed temperature), where the Lorenz ratio deviates from one. First, we have used standard solid state book-based non-fluid expressions of Lorenz ratio for the 3D-NR case, which are also applied for 2D-NR, 3D-G, and 2D-G cases. For all cases, the Lorenz ratio remains constant in the higher values of normalized Fermi energy. This fact is well agreement with the universal validity of the WF law in metal systems, which can be described by the 3D-NR case of the electron gas having larger Fermi energy (2102-10 eV) or normalized Fermi energy 400-2000 at $60^{\circ}$ K. Depending on the definition of specific heat at constant volume and Fermi energy, we will get two possible saturated values of Lorenz ratios - 1 and 2, where former is experimentally observed (average) value but one may mark the saturating domain as the WF law obeying domain. If one identifies two domains - less and greater than $10$ value of normalized Fermi energy (i.e. $\frac{\epsilon_{F}}{T}<10$ and $\frac{\epsilon_{F}}{T}>10$), then the former domain of all cases show the deviation of WF law. This tuning of normalized Fermi energy is conveniently possible for 2D-G system only via tuning the electron doping. Recently measured Lorenz ratio by Crossno et al. [crossno2016observation] indicates the violation of the WF law in the lower normalized Fermi energy domain, which is popularly called as Dirac fluid (DF) domain. Using standard solid state or non-fluid theory for 2D-G system, we also find WF law violation in the lower normalized Fermi energy domain. However, non-fluid theory gets less than one values of Lorenz ratio, while experiment gets its greater than one values. It demands that we may have to transist towards a fluid-type from the standard non-fluid expressions of the Lorenz ratio. So far, to the best of our knowledge, this straightforward thermodynamical phase space analysis of Lorenz ratio, based on non-fluid theory is missing in the literature. Present work is fulfilling this gap, which may be considered one of the important contents of present article. A next level content of present article is its fluid based calculation,whose outcomes are summarized in the next paragraph.
@@
Towards the fluid description, we have approached in two steps. In the first step, we have used Boltzmann transport equation for 2D-G system, where local ther- malization concept is mainly implimented for fluid de- scription. If we compare the NF and F expressions of Lorenz ratios, then we can find that only via specific heat, Fermi integral functions are entered into the NF expression of Lorenz ratio. On the other hand, F expressions of electric and thermal conductivity and hence Lorenz ratio carry additional Fermi integral functions, which are com- pletely appeared because of local thermalization aspect of fluid description. During this first step attempt of fluid description, we get a little deviated outcomes of Lorenz ratio with respect to NF results. Sommerfeld limit of F-based Lorenz ratios give a divergent outcomes instead of constant values (it was found in its NF-based expression). This divergent trend of Lorenz ratio in lower nor- malized Fermi energy is experimentally observed, so this first step fluid description results may be considered as a preliminary hints of NF to F transition in the Dirac fluid domain or lower Fermi energy domain. Observing the quantitative mismatch between experimental data and first step fluid theory, we have proceeded to second step extension of fluid theory. Here, we have gone through the replacement of specific heat by enthalpy per net electron (divided by temperature). This replacement is based on the intuition for transition from quasi-particle concept to fluid concept. In NF picture, electron average energy is the quasi-particle concept and its temperature deriva- tive gives specific heat concept. Whereas, enthalpy per net electron is the fluid concept for electron-hole plasma in Dirac fluid domain. Interestingly, we notice that this conceptual transformation and replacement provide an enhancement of Lorenz ratio in Dirac fluid domain and their theoretical values are quite close to experimental data points.
V Conclusion
@@
Present work has explored the WF law violation of electrons in graphene in terms of a non-fluid to fluid transition framework. After developing a step by step NF to F frameworks, we have identified four regions along the axis of normalized Fermi energy $\frac{\epsilon_{F}}{T}$ at $T=60^{\circ}$ K:
@@
  • $\frac{\epsilon_{F}}{T}>10$: Lorenz ratio is one and it can be called hardcore NF Ohmic domain. Sommerfeld limit of our NF expression can explain it easily.
@@
  • Within $2.5<\frac{\epsilon_{F}}{T}<10$: Lorenz raio become less than one and it can be called NF but non-ohmic type domain. Our NF expression of Lorenz ratio and experimental data both hint this domain.
@@
  • $0.7<\frac{\epsilon_{F}}{T}<2.5$: It is the transition domain from NF to F by showing less than one to greater than one values of Lorenz ratio.
@@
  • $\frac{\epsilon_{F}}{T}<0.7$: Here, Lorenz ratio becomes greater than one and it can be called Fluid domain as enhancing trend in fluid theory and experimental data towards lower doping range is noticed.
@@
At the end, for final quantitative data matching, we have demonstrated a sandwich estimation of F and NF results. However, present article intend to draw the attention of reader on the qualitative message of NF to F transition. This association of enhancement Lorenz ratio observation in the Dirac fluid domain with the NF to F transition is agreement with earlier theoretical works.
@@ Appendix A FERMI-DIRAC DISTRIBUTION FUNCTION
$\frac{\partial f_{0}}{\partial\epsilon}=\frac{\beta e^{\beta\left(\epsilon-\epsilon_{F}\right)}}{\left(e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1\right)^{2}}=\frac{\partial}{\partial\epsilon_{F}}\left(\frac{1}{e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1}\right).$ (45)
@@ Appendix B FERMI-DIRAC FUNCTION
    Case.2. When $A$ is too much small $\left(\epsilon_{F}<<k_{B}T\right)$, then the function becomes simplified as AA
@@ Appendix C ELECTRONIC SPECIFIC HEAT
$N=\int_{0}^{\infty}D\left(\epsilon\right)d\epsilon f_{0}\left(\epsilon\right).$ (55)
@@
$n=\frac{N}{S}=\frac{g}{2\pi\hbar^{2}v_{F}^{2}}\int_{0}^{\infty}\frac{\epsilon}{A^{-1}e^{\beta\epsilon}+1}d\epsilon.$ (56)
@@
$U=\int_{0}^{\infty}D\left(\epsilon\right)d\epsilon f_{0}\left(\epsilon\right)\epsilon.$ (59)
@@
$\prescript{2D}{G1}{[C_{V}]_{e1}}=\Bigg[\frac{\partial}{\partial T}\Big(\frac{U}{N}\Big)\Bigg]_{S,\ \epsilon_{F}}.$ (63)
@@
$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg[\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}+T\frac{\partial}{\partial T}\left(\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\right)\Bigg].$ (64)
@@
$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg]~.$ (66)
@@
$\prescript{2D}{G2}{[C_{V}]_{e2}}=\frac{1}{N}\frac{\partial U}{\partial T}\Bigg|_{S,\ \epsilon_{F}}.$ (67)
@@
$\prescript{2D}{G2}{[C_{V}]_{e2}}=2k_{B}\Bigg[3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg].$ (70)
@@
$\prescript{2D}{NR1}{[C_{V}]_{e1}}=k_{B}\Bigg[2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg],$ (71)
@@
$\prescript{2D}{NR2}{[C_{V}]_{e2}}=k_{B}\Bigg[2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg].$ (72)
@@
$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{3}{2}k_{B}\Bigg[\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg],$ (73)
$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3}{2}k_{B}\Bigg[\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg],$ (74)
$\prescript{3D}{G1}{[C_{V}]_{e1}}=3k_{B}\Bigg[4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg],$ (75)
$\prescript{3D}{G2}{[C_{V}]_{e2}}=3k_{B}\Bigg[4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg].$ (76)
Appendix D NON-FLUID DESCRIPTION of OTHER SYSTEMS
3D-NR: The well-known Drude’s formula for the electrical conductivity of nonrelativistic electrons in 3D solids (metals) is
$\sigma_{NR}^{3D}=\frac{ne^{2}\tau_{c}}{m}=\frac{ne^{2}\lambda}{mv_{F}},$ (77)
and the thermal conductivity is
$\kappa_{NR}^{3D}=\frac{1}{3}nv_{F}\lambda\,\prescript{3D}{NR}{[C_{V}]_{e}},$ (78)
where, $n$ is the number (electron) density, $\tau_{c}$ is the relaxation time, $m$ is electron mass, $\lambda=v_{F}\tau_{c}$ is the mean free path, and $C_{V}]_{e}$ is the electronic specific heat per particle. After taking the ratio of thermal and electrical conductivity, we get
where $\epsilon_{F}=\mu=\frac{1}{2}mv_{F}^{2}$ is the Fermi energy.
Here, in Eq. (79), $C_{V}]_{e}$ is the main important thing to calculate the conductivity ratio. We will use the two definitions of $C_{V}]_{e}$:
  1. 1.
    $C_{V}]_{e1}=\frac{\partial}{\partial T}\Big(\frac{U}{N}\Big)\Bigg|_{V,\ \epsilon_{F}}$,
  2. 2.
    $C_{V}]=\frac{\partial}{\partial T}\Big(\frac{U}{N}\Big)\Bigg|_{V,\ \epsilon_{F}}$,
  3. 3.
    $C_{V}]_{e2}=\frac{1}{N}\frac{\partial U}{\partial T}\Bigg|_{V,\ \epsilon_{F}}$.
Here, for subsequent simplicity, we will use the notations like NR1 and G1 concerning $C_{V}]_{e1}$ and NR2 and G2 dealing with $C_{V}]_{e2}$. The former definition prescribes taking $T$ derivative of internal energy per particle $u=U/N$, while the latter definition says to take $T$ derivative of internal energy $U$ and then normalized by $N$. The detailed calculation can be seen in Appendix C. Let us first put the former specific heat,
$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{3}{2}k_{B}\Bigg[\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg],$ (80)
in Eq. (79), the Lorenz ratio for 3D-NR will be
Here,
$f_{\nu}(A)=\frac{1}{\Gamma(\nu)}\int_{0}^{\infty}\frac{x^{\nu-1}}{A^{-1}e^{x}+1}dx,$ (82)
is the standard Fermi integral function (see details in Appendix B) with $A=e^{\epsilon_{F}/k_{B}T}$. Using Sommerfeld lemma, in the limit of $\epsilon_{F}/k_{B}T=lnA>>1$, the electronic specific heat becomes [ashcroft2022solid]
$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{\pi^{2}}{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (83)
and then Eq. (81) becomes
$L_{NR1}^{3D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (84)
which is the so-called Wiedemann-Franz law for metals.
Next, using another definition of $C_{V}]_{e2}$ for the 3D-NR case, whose general expression will be (see Appendix C)
$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3}{2}k_{B}\Bigg[\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg]~,$ (85)
we will get the Lorenz ratio as
$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3\pi^{2}}{4}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (87)
and then the Eq. (86) becomes
$L_{NR2}^{3D}=\frac{\pi^{2}}{2}\left(\frac{k_{B}}{e}\right)^{2}=1.5L_{0}.$ (88)
Reader can see that Eq. (84) and (88), both are independent of $T$ , $\epsilon_{F}$ and are made by universal constants like $k_{B}$ and $e$. So, the domain of $T-\epsilon_{F}$ plane, where $\frac{L}{L_{0}}$ or $L$ becomes independent of $T$ , $\epsilon_{F}$ , can be considered as Wiedemann-Franz law obeying domain and remaining will be Wiedemann-Franz law violating domain. Although, in terms of $_{0} , LL$ is $1.5$ factor times larger for Eq. (88).
and the thermal conductivity,
$\kappa_{G}^{3D}=\frac{1}{3}nv_{F}\lambda\prescript{3D}{G}{[C_{V}]_{e}},$ (90)
will form the ratio
$\prescript{3D}{G1}{[C_{V}]_{e1}}=3k_{B}\Bigg[4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg]$ (92)
and
$\prescript{3D}{G2}{[C_{V}]_{e2}}=3k_{B}\Bigg[4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg],$ (93)
in Eq. (91), we get the Lorenz ratios
and
$\prescript{3D}{G1}{[C_{V}]_{e1}}=\pi^{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (96)
and
$\prescript{3D}{G2}{[C_{V}]_{e2}}=3\pi^{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (97)
and so, Eqs. (94) and (95) become
$L_{G1}^{3D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (98)
and
$L_{G2}^{3D}=\pi^{2}\left(\frac{k_{B}}{e}\right)^{2}=3L_{0},$ (99)
respectively.
$\prescript{2D}{NR1}{[C_{V}]_{e1}}=k_{B}\Bigg[2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg],$ (101)
and
$\prescript{2D}{NR2}{[C_{V}]_{e2}}=k_{B}\Bigg[2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg],$ (102)
in Eq. (100), we get the Lorenz ratios
and
$L_{NR1}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (105)
$L_{NR2}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0}.$ (106)
Appendix E APPROACH TOWARDS THE FLUID DESCRIPTION OF OTHER SYSTEMS
2D-NR After doing the same calculation as the 2D G of towards fluid system for a 2-dimensional nonrelativistic system, the final general expressions of the electrical and two possible thermal conductivities and their corresponding Lorenz ratios $L_{NR}^{2D}/L_{0}$ are given by as below:
$f_{\frac{5}{2}}\left(A\right)T^{\frac{5}{2}},$ (118)
$f_{\frac{5}{2}}\left(A\right)T^{\frac{5}{2}},$ (119)